OpenStudy (anonymous):

solve : ln5x+ln(x+1)=0

OpenStudy (anonymous):

so far I've done lin5x(x+1)= 0 ln5x^2+5x=0 ln5(x^2+x)=0

OpenStudy (anonymous):

and 0= ln1

OpenStudy (anonymous):

where do I go from there?

OpenStudy (anonymous):

then remove log and you'll have 5x^2+5x-1=0

OpenStudy (anonymous):

ln5x+ln(x+1)=0 ln(5x^2+5x) =0 5x^2 +5x=1 solve it through completing square method you will be getting the answer

OpenStudy (anonymous):

where does the 1 come from?

OpenStudy (anonymous):

0 is equal to ln 1

OpenStudy (anonymous):

just factor out the ln and devide both sides by it...u end up with 5x+(x+1)

OpenStudy (ash2326):

ln 5x+ln x+1=0 ln(5x*(x+1))=0 taking anti log on both sides e^0=1 so 5x(x+1)=1 5x^2+5x-1=0 solving using quadratic formula we get \[x=(-5+\sqrt(5^2-4*5*-1))/10\] or \[x=(-5+\sqrt(45))/10\] \[x=(-5+3\sqrt(5))/10\] and \[x=(-5-3\sqrt(5))/10\]

OpenStudy (anonymous):

@Jason Good heavens don't say divide by the log sign.... *restarts heart*

OpenStudy (anonymous):

thank you everyone for helping me! @uselessmass people are being nice and helping me out, just like you...you added to the post!

OpenStudy (anonymous):

A) He says as he adds another one... :) and B) that should be \[(1+x)(5x) = 1, \space \text{not} \space 0\]

OpenStudy (anonymous):

still kinda confused but it's getting there!

OpenStudy (anonymous):

so I got 5x^2+5x-1 and now i use the quadratic formula right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay so I got final answer as [-5-3\sqrt{5}/10\]

OpenStudy (anonymous):

\[-5-3\sqrt{5}/10\]